OpenStudy (kikilandry):

3^2 + (7 − 2) ⋅ 4 − 6 over 3.

alones (alones):

Is this what you mean? @kikilandry \[\frac{ 3^2 + (7 − 2) ⋅ 4 − 6 }{ 3 }\]

OpenStudy (michaelschmidt):

i didnt post this

OpenStudy (kikilandry):

\[3^2+ (7-2)\times 4- \frac{ 6 }{ 3}\]

OpenStudy (proheiper):

lmao always blaming michael for stuff

alones (alones):

@kikilandry so you got your own answer, or you need help with it?

OpenStudy (kikilandry):

i need help with it

alones (alones):

Sure. \[9+(7−2)(4)-\frac{ 6 }{ 3 }\] \[9+(5)(4)−\frac{ 6 }{ 3 }\] \[9+20−\frac{ 6 }{ 3 }\] \[29−\frac{ 6 }{ 3 }\] \[29−2 =27\]

OpenStudy (kikilandry):

thank you so much!

alones (alones):

You welcome hun